### v1

# OMP4Py: a pure Python implementation of OpenMP

Abstract
Python demonstrates lower performance in comparison to traditional high performance computing (HPC) languages such as C, C++, and Fortran. This performance gap is largely due to Python’s interpreted nature and the Global Interpreter Lock (GIL), which hampers multithreading efficiency. However, the latest version of Python includes the necessary changes to make the interpreter thread-safe, allowing Python code to run without the GIL. This important update will enable users to fully exploit multithreading parallelism in Python. In order to facilitate that task, this paper introduces OMP4Py, the first pure Python implementation of OpenMP. We demonstrate that it is possible to bring OpenMP’s familiar directive-based parallelization paradigm to Python, allowing developers to write parallel code with the same level of control and flexibility as in C, C++, or Fortran. The experimental evaluation shows that OMP4Py significantly impacts the performance of various types of applications, although the current threading limitations of Python’s interpreter (v3.13) reduce its effectiveness for numerical applications.
keywords:
OpenMP , Python , Parallelism , Multithreading , Scalability
journal: Future Generation Computer Systemst1t1footnotetext: This work was supported by MICINN [PLEC2021-007662, PID2022-137061OB-C22]; Xunta de Galicia [ED431G 2019/04, ED431F 2020/08, ED431C 2022/16]; and European Regional Development Fund (ERDF).
1 Introduction
Lately, Python has become the most popular programming language [1]. Its ease of use has led to its widespread adoption across many scientific domains, from data analysis to machine learning. Despite its advantages in these areas, Python significantly lags behind traditional low-level HPC (High Performance Computing) languages like C/C++ and Fortran when it comes to achieving high performance. This performance gap can be attributed to two primary factors. First, Python’s interpreted nature introduces overhead that impacts execution speed compared to the compiled code of languages such as C and Fortran. Second, Python’s Global Interpreter Lock (GIL) limits its ability to fully exploit multithreading, hindering its scalability in parallel computing tasks. These limitations present challenges to utilizing Python in HPC environments, where speed and efficiency are crucial.
Various efforts have been made to improve support for multithreading parallelism in Python, but important limitations remain [2]. However, a more definite solution is expected with the latest Python interpreter, whose final release was in October 2024. This version includes the necessary changes to make the interpreter thread-safe, allowing Python code to run without the GIL.
Based on this important advancement, this paper introduces OMP4Py [^1] It is publicly available at https://github.com/citiususc/omp4py, the first native Python implementation of OpenMP [3], a widely recognized standard programming model for exploiting multithreading parallelism in HPC. OMP4Py integrates OpenMP functionalities into Python through two main mechanisms. First, it adapts OpenMP’s directive-based approach, allowing Python users to embed parallel constructs directly into their source code. These transformer directives enable parallel execution by instructing OMP4Py to modify the Python code accordingly. Second, OMP4Py includes a set of runtime library functions that closely mirror those provided by OpenMP. These functions allow users to manage parallel execution parameters, such as the number of threads and scheduling policies, giving them flexibility to fine-tune the parallel behavior of their Python programs. In this way, we prove that OpenMP’s directive-based parallelization model can be seamlessly integrated into Python, giving developers the same level of control in writing parallel code as they would have in C, C++, or Fortran.
A thorough experimental evaluation was carried out in the paper, demonstrating that OMP4Py has significant potential for hybrid parallelism in combination with mpi4py [4] and for non-numerical workloads. However, the current threading limitations of Python’s interpreter (v3.13) hinder its scalability for numerical applications.
The paper is structured as follows: Section 2 provides some background and summarizes previous research in the field. Section 3 introduces OMP4Py, detailing its design, features, and implementation. Section 4 presents the experimental results, discussing various experiments conducted to evaluate OMP4Py’s performance across different types of applications. Finally, Section 5 summarizes the key findings and proposes directions for future research..
2 Background & Related work
2.1 OpenMP
OpenMP [3] is a parallel programming model designed for shared-memory computer systems, aiming to simplify the exploitation of inherent concurrency in many algorithms. OpenMP uses the fork-join model of parallel execution. In this model, the program starts with a single master thread. At parallel regions, the master thread creates multiple parallel threads that concurrently execute the assigned tasks. Once the parallel tasks are completed, the threads join back into the master thread, which continues executing the program sequentially. Through compiler directives, programmers can instruct the compiler to generate multithreaded code at a higher level. The OpenMP API is compatible with C/C++ and Fortran. Unlike lower-level approaches such as pthreads in the POSIX library, where the programmer manually manages thread creation and task assignment, OpenMP abstracts these complexities, providing a higher-level programming model. For this reason, OpenMP is considered by the HPC community as the standard for exploiting the multithreading capabilities of current multi-core CPUs.
The OpenMP API standard specification [^2] https://www.openmp.org/specifications [online, accessed November 5, 2024] started in 1997 (version 1.0), and it continues to evolve, with new constructs and features being added over time. The latest release, version 6.0, was recently published in November 2024. However, most OpenMP programmers typically use only a subset of the OpenMP 3.0 specification released in 2008. This subset, referred to as the OpenMP Common Core [5], comprises the 21 most commonly utilized elements of OpenMP. Here are some of the main constructs, clauses and functions included in the OpenMP Common Core:
  1. 1.
    Parallel regions:
    • (a)
      #pragma omp parallel - Defines a parallel region, which is a block of code executed by multiple threads.
  2. 2.
    Work sharing constructs:
    • (a)
      #pragma omp for - Distributes loop iterations among threads in a parallel region. There also exists a combined construct equivalent to a parallel construct followed by a for (#pragma omp parallel for).
    • (b)
      #pragma omp sections - Divides work into separate sections that can be executed in parallel.
    • (c)
      #pragma omp single - Specifies a block of code that should be executed by only one thread.
  3. 3.
    Tasking:
    • (a)
      #pragma omp task - Creates an explicit task for deferred execution within the construct.
    • (b)
      #pragma omp taskwait - Ensures that all child tasks created within the current task are completed before the execution of the program continues.
  4. 4.
    Synchronization constructs:
    • (a)
      #pragma omp barrier - All threads must reach a barrier before any can continue.
    • (b)
      #pragma omp critical - Ensures that only one thread at a time executes a block of code.
  5. 5.
    Scheduling and other clauses:
    • (a)
      schedule(static [,chunk]) - Distributes the iterations of a loop in contiguous blocks, with each thread receiving a block of iterations of size chunk.
    • (b)
      schedule(dynamic [,chunk]) - Dynamically assigns chunks of iterations to threads at runtime.
    • (c)
      nowait -It is used to remove the implicit barrier at the end of certain constructs such as for, sections and single.
  6. 6.
    Data environment clauses:
    • (a)
      private(list) - Specifies that each thread should have its own private copy of the variables in the list.
    • (b)
      firstprivate(list) - Similar to private, but the initial value of the variable is copied from the master thread.
    • (c)
      lastprivate(list) - Ensures that the value of a private variable from the last iteration is copied back to the original variable.
    • (d)
      shared(list) - Specifies that variables in the list should be shared among all threads.
    • (e)
      reduction(op:list) - Performs a reduction operation on variables across all threads in a parallel region.
  7. 7.
    Functions:
    • (a)
      omp_set_num_threads(int) - Sets the number of threads that will be used in the next parallel region.
    • (b)
      omp_get_thread_num() - Returns the unique thread number of the calling thread within its team.
    • (c)
      omp_get_wtime() - Returns the elapsed time since some point in the past. It is used to measure the execution time of a segment of code.
2.2 Limitations of Python for multithreading parallelism
As commented previously, Python has emerged as the most popular programming language in recent years [1], renowned for its simplicity, productivity, and readability. As a result, we can find Python applications in practically all scientific areas. However, when the main goal is obtaining high performance, Python falls far behind traditional low-level HPC languages such as C and Fortran. There are two main reasons for that behavior.
First, Python is an interpreted programming language, which causes an important overhead due to the need for real-time translation of source code into machine code during runtime. This problem can be mitigated using Just-In-Time (JIT) compilers like Numba [6], which translates Python functions to optimized machine code at runtime using the LLVM compiler. This approach aims to achieve performance levels comparable to those of C or Fortran. However, Numba is optimized for numerical computations, so code that involves extensive string manipulation, complex data structures, or I/O operations may not see significant performance gains. This is the case, for example, of purely Pythonic code that relies heavily on built-in data structures like lists and dictionaries.
The second issue is related to how Python handles multithreading due to the existence of the Global Interpreter Lock (GIL). The GIL is a locking mechanism that protects access to Python objects, preventing multiple threads from executing Python code simultaneously. The GIL was originally introduced to simplify thread management and protect against race conditions and memory corruption, making it easier for developers to write concurrent code safely. However, since the GIL allows only one thread to execute Python code at a time, multithreading in Python is unsuitable for CPU-bound tasks where the performance gain from parallel execution is significant. In this way, the GIL becomes the most important obstacle to take advantage of parallelism using multi-core CPUs efficiently in Python. Note that, on the other hand, I/O-bound tasks, such as network requests or file operations, can benefit from multithreading in Python. In these cases, the GIL is released when a thread performs I/O operations, allowing other threads to execute Python code.
2.3 Advances towards a multithreading-friendly Python
Python currently offers several methods to enable parallelism, but these techniques have significant limitations [2]. For instance, the multiprocessing library allows programs to create and interact with subprocesses. This facilitates parallelism because each subprocess has its own Python interpreter, meaning there is one GIL per process. Some examples of applications and libraries that use the multiprocessing package to parallelize tasks are the deep learning frameworks PyTorch [7] and TensorFlow [8], and the HPC-Big Data framework IgnisHPC [9]. However, multiprocessing has some important drawbacks. Communication between processes is limited, as objects typically need to be serialized or copied to shared memory. This introduces overhead and complicates building APIs on top of multiprocessing. In addition, starting a subprocess is more expensive than starting a thread. Finally, many C and C++ libraries support multithreading but do not support access or usage across multiple processes.
An additional solution to take advantage of parallelism in Python (multithreading in this case) is based on the observation that functions implemented in C may use multiple threads internally. For instance, Intel’s NumPy distribution, among others, employs this approach to internally parallelize individual operations. This works well when the basic operations are large enough to be parallelized efficiently, but not when there are many small operations or when the operations depend on some Python code. Note that invoking Python from C requires obtaining the GIL, which means even small pieces of Python code can prevent scaling.
Another interesting alternative to bring parallel multithreading to Python is PyOMP [10, 11], which is a prototype system with support for OpenMP. As noted earlier, OpenMP API tells the compiler how to generate multithreaded code. In the case of PyOMP, Numba acts as the compiler, transforming Python code into LLVM, thereby bypassing the GIL. Note that the Numba compiler works with NumPy arrays, which must be used for any arrays inside a PyOMP function. However, as stated previously, Numba, and consequently PyOMP, struggle when calling functions or interacting with Python objects. Due to the lack of compiler directives in Python, PyOMP uses the with statement instead. In this way, for example, to create a parallel region uses: with openmp("parallel"):, while the equivalent directive in C/C++ API would be: #pragma omp parallel. On the other hand, PyOMP supports 90% of the OpenMP Common Core missing nowait and the dynamic schedule (see Section 2.1).
As we have already discussed, the main factor that severely limits the concurrency of multithreaded Python code is the GIL. Despite several efforts over the years to remove the GIL [12, 13], none have been considered for inclusion in the Python interpreter until recently [2]. As a result, Python 3.13, whose final release was in October 2024, includes a build configuration flag (--disable-gil) to allow it to run Python code without the Global Interpreter Lock and with the necessary changes needed to make the interpreter thread-safe. This will initiate a long process toward making the disabling of the GIL the default option in the Python interpreter. Note that a basic JIT compiler was also added to Python 3.13 [14], aiming to reduce the performance distance with compiled languages such as C and C++.
3 OMP4Py
OMP4Py is a novel implementation of the OpenMP standard, designed specifically for Python. It currently supports the complete specifications of version 3.0. This tool was developed with adherence to the official OpenMP documentation [^3] https://www.openmp.org/wp-content/uploads/spec30.pdf [online, accessed November 5, 2024], which outlines the necessary compiler directives, runtime library functions, and environment variables required for creating shared memory parallel programs using threads. The OpenMP standard traditionally supports C, C++, and Fortran, which are low-level, compiled languages that use preprocessor directives to manage parallelism. These directives are not part of the core language syntax but are preprocessor instructions that modify the source code prior to the compilation process. In C/C++, these directives are introduced using #pragma, and in Fortran, they are marked by the sentinel !\$. The main goal of OMP4py is to bring the familiar parallelization paradigm of OpenMP to Python, allowing Python developers to write parallel code with the same level of control and flexibility as in C, C++, or Fortran. This tool aims to port the OpenMP model, with all code execution handled natively using Python threads. This ensures integration with Python’s libraries, enabling multi-threaded performance directly within Python’s ecosystem.
OMP4Py integrates the core functionalities of OpenMP into Python in the following ways:
  • 1.
    Transformer directives: Adapting OpenMP’s directive-based approach, OMP4Py allows Python users to embed parallel constructs directly into their source code. These directives instruct the OMP4Py to transform the Python code for parallel execution.
  • 2.
    Runtime library functions: OMP4Py includes a set of runtime library functions that mirror those provided by OpenMP. These functions manage parallel execution parameters, such as the number of threads and scheduling policies, among others. This feature provides users the flexibility to fine-tune the parallel behavior of their Python programs.
Since Python lacks a preprocessor, we needed to integrate OpenMP directives directly into the Python language while adhering to Python’s best practices and idioms. To achieve this, we defined a function omp that operates similarly to OpenMP directives in C/C++, maintaining the same syntax and functionality. The function itself has no effect when executed; it serves solely as a container for the OpenMP directives. In this way, preprocessor directives in C/C++, such as:
#pragma omp parallel num_threads(2)
would be integrated into Python using OMP4Py as:
with omp("parallel num_threads(2)"):
Note that when a OpenMP directive must be used within structured blocks, the omp function is used together as part of a with block (similar to PyOMP syntax [11]); otherwise, it is used as a standalone function call. Finally, by themselves, calls to the omp function, as mentioned earlier, have no effect. Therefore, we must instruct the Python interpreter to restructure the code according to the content of each OpenMP directive before executing it. To accomplish this task, we employ Python decorators. A decorator can be applied to a function or class to change its behaviour in a very elegant and intuitive way. So, we must decorate a function or class containing the OpenMP directives with the @omp decorator.
1 from omp4py import *
2 import random
3
4 @omp
5 def pi(num_points):
6 count = 0
7 with omp("parallel for reduction(+:count)"):
8 for i in range(num_points):
9 x = random.random()
10 y = random.random()
11 if x * x + y * y <= 1.0:
12 count += 1
13 pi = count / num_points
14 return pi
15
16 print(pi(10000000))
Figure 1: Example of a Monte Carlo method for $\pi$ calculation using OMP4Py.
Figure 1 shows an example of Python code for the parallel calculation of $\pi$ using OMP4Py. First, in line 4, the pi function is decorated with @omp, indicating that it contains OpenMP directives that need to be processed. Next, in line 7, a parallel region is started using omp("parallel for reduction(+:count)"). This statement instructs OMP4Py to parallelize the following for loop, where each thread contributes to the reduction operation on count. Finally, line 14 returns the computed value of $\pi$.
To achieve parallel code generation in Python, we must consider that Python is an interpreted language. Unlike lower-level languages defined in the OpenMP standard, there is no compilation process where we can apply the transformations needed to produce parallel code. Instead, following Python’s philosophy, OpenMP code is generated by the interpreter itself at the moment the module containing the user’s code is loaded. For a Python module to begin execution, all global definitions must be loaded, which can include the importation of other modules, global variables, and, relevant to our case, the declaration of functions or classes. When we apply a decorator to a function or class, the interpreter will invoke this decorator with the function or class as arguments right after it has finished loading it. The result of executing the decorator will replace the original function or class, and once the module is fully loaded, the user’s code will only interact with the decorated version.
The omp decorator aims to replace all directives within the source code of a function or class and generate a new version with parallel behavior. The first step is to obtain information about the object representing the decorated function or class. For this, we use the inspect module, which allows us to retrieve the source code of the object. Once we have the source code, we generate an abstract syntax tree (AST) using Python’s ast module. The AST provides a simple and transformable representation of the source code. The transformation process involves an in-order traversal of the AST. Each time a directive is encountered, it is parsed and checked for errors before applying the transformations to the AST. If any errors are found, the interpreter will abort with a SyntaxError, just as it would when encountering invalid syntax in Python code. After all directives have been processed, the resulting tree is transformed into object code using the compile function and loaded into the interpreter with exec. Finally, the decorator returns the new function or class with the parallel code, replacing the original.
3.1 Parallel Directive
In OpenMP, the parallel directive is essential for creating parallel programs. This construct allows a specific segment of code to execute simultaneously across multiple threads. When a code block is preceded by this construct, the program creates a team of threads, and each thread, including the original one, executes a copy of the code block concurrently. Multiple parallel directives can be nested, enabling each thread to create new teams of threads that work independently. This feature is known as nested parallelism and must be enabled with omp_set_nested. By definition in the standard, every OpenMP program runs within an implicit single-threaded parallel construct. This allows API functions like omp_get_num_threads or omp_get_thread_num, among others, to operate independently of where they are called in the code. If nested parallelism is disabled, only an explicit parallel execution is allowed; subsequent parallel regions are ignored and executed sequentially by the threads that encounter them.
Variables defined within a parallel block are local to each thread, meaning that each thread has its own copy. However, previously defined variables can be either private or shared, according to the user’s preference. Shared variables maintain the same value across all threads. Any modifications made by one thread are visible to the others, and this shared value persists even after the parallel block ends. By default, all variables defined before a parallel block are shared. While users can explicitly mark variables as shared, they only need to specify which variables should be private. Declaring a variable private makes it behave as if it was created inside the parallel block, causing its initial value to be lost when the block ends. To retain the value of a variable, users can employ firstprivate, which initializes each thread’s local copy to the variable’s previous value.
1from omp4py import *
2
3@omp
4def f(a):
5 b = "1"
6 c = -1
7 d = [1, 2]
8 e = True
9 with omp("parallel shared(b) private(c) firstprivate(d) num_threads(4)"):
10 f = omp_get_thread_num()
11 a = 1
12 c = f
13 d.append(3)
14 print(b, c, d, f)
15
16 print(a, c)
1from omp4py import *
2
3def f(a):
4 b = "1"
5 c = -1
6 d = [1, 2]
7 e = True
8 def _omp_parallel_1():
9 nonlocal a, b
10 _omp_c_2 = None
11 _omp_d_3 = _omp_copy(d)
12 f = omp_get_thread_num()
13 a = 1
14 _omp_c_2 = f
15 _omp_d_3.append(f)
16 print(b, _omp_c_2, _omp_d_3, f)
17
18 _omp_parallel_run(_omp_parallel_1, num_threads=4)
19
20 print(a, c)
Figure 2: Example of the parallel directive: user code (top) and its corresponding translation by OMP4Py (bottom).
The implementation of the parallel directive in Python presents two main challenges. First, the code that is meant to run in parallel across different threads must be placed inside a function. Second, it is necessary to manage variables used both inside and outside the code block to determine whether they should be shared or private for each thread. Figure 2 shows an example of a user-defined parallel block (top) and the corresponding code generated by OMP4Py for its parallel execution (bottom). It is important to note that, in code generation examples, the prefix _omp_ will be used for all internal OMP4Py symbols, with dynamically created symbols including a number to avoid collisions. The example shows a simple function with a parallel directive and six representative variables (’a’ to ’f’) with different types. The first change to observe is that the decorator (line 3, top code) and the function containing the directive (line 9, top code) have been removed. Once the code is transformed, the decorator is no longer needed and must be removed to prevent multiple processing. The parallel directive (line 9, top code) has been transformed into a function (line 8, bottom code) that will be called by different threads. The _omp_parallel_run function (line 18, bottom code) is an internal OMP4Py function responsible for initializing and launching threads using Python’s Thread class. Since the user has specified a number of threads with num_threads, we also need to pass this argument. Note that the existing thread, or master thread, is an execution thread as well, so only $num\_threads-1$ additional threads are created.
The next step is to deal with local variables. The parallel block (lines 9-16, bottom code) is now within a nested function, a term for functions that are defined inside another function. In Python, nested functions can access variables from the outer function. First, variables a and b are shared; in this case, we implicitly use the nonlocal keyword (line 9, bottom code) to indicate that any new assignment modifies the variable in the outer function rather than creating a new one inside the nested function. Additionally, variables c and d are declared private, so we create two new variables and replace all their uses within parallel block code. Variable c has no initial value (line 10, bottom code), while d is initialized using _omp_copy, which creates a shadow copy of the variable (line 11, bottom code). Finally, variable e is not used within the code, and f is a local variable, so they are ignored. Consequently, we observe that the internal print statement (line 16, bottom code) will show the same shared value of b in each thread, while the other values will differ for each thread. However, the external print statement (line 20, bottom code) will display the modified value of a from within the parallel block (line 13, bottom code) but will retain the initial value of c (line 5, bottom code).
3.2 Worksharing Constructs
The distribution of work among threads is the base of OpenMP’s functionality. A worksharing construct divides execution regions among the team of threads created in the most recent parallel directive. Common worksharing constructs include for loops, which distribute loop iterations among threads, sections, which divide distinct code blocks among threads, and single, which ensures that a specific section of code is executed by only one thread. Worksharing regions are executed as soon as the first thread encounters them. While the first thread can begin executing its part of the work immediately, it must remain in the worksharing region until all threads in the team reach the same point. This is the default behavior for all worksharing constructs, although users can modify it if necessary using the nowait clause. If there is only a single thread, no work distribution occurs, and the worksharing regions are executed sequentially.
3.2.1 For
The for directive is used to parallelize loops, allowing iterations to be distributed across multiple threads. Iterations are grouped into chunks and assigned to each thread according to a scheduling policy. OpenMP defines three scheduling policies: static, dynamic, and guided. The static policy assigns chunks to threads in a round-robin fashion, the dynamic policy assigns chunks as they are requested by threads using a shared variable index, and the guided policy is similar to dynamic but the chunk size decreases dynamically with each assignment. Additionally, auto allows the compiler to choose the scheduling policy, while runtime defers the decision to runtime, where it is determined by an environment variable.
The first challenge to implement the for directive in Python is the absence of the traditional for loop. In Python, all for loops are constructed as foreach loops that iterate over a collection of elements. Consequently, to achieve the equivalent of a classic for loop, one must use a foreach loop over a range of values generated by the built-in range function. For example, a simple C-style loop like for(int i = 0; i < 10; i++) is equivalent to for i in range(0, 10, 1) in Python, where the initial value, final value, and step are specified as arguments.
1from omp4py import *
2
3@omp
4def f():
5 xs = [0] * 20
6 with omp("parallel"):
7 with omp("for schedule(static, 2)"):
8 for i in range(len(xs)):
9 xs[i] = i
10 print(xs)
1from omp4py import *
2
3def f():
4 xs = [0] * 20
5 def _omp_parallel_1():
6 nonlocal xs
7 for i in _omp_range(0, len(xs), 1, schedule=’static’, chunks=2):
8 xs[i] = i
9
10 _omp_parallel_run(_omp_parallel_1)
11
12 print(xs)
Figure 3: Example of the for directive: user code (top) and its corresponding translation by OMP4Py (bottom).
Figure 3 shows an example of a loop parallelized with the for directive (top) and its internal implementation generated by OMP4Py (bottom). First, we need to note that the for directive must be within a parallel directive to distribute the work (lines 6-7, top code). Unlike the parallel directive, the for directive must be the unique element within the with block; no other statements are allowed inside the block except within the loop itself. The implementation of this directive is simple: the range function (line 8, top code) is replaced by _omp_range (line 7, bottom code), and the scheduling and chunk values are passed as keyword arguments to the function. The _omp_range function is always generated with the three positional arguments of range, even if the user does not explicitly specify them. The function, implemented in the OMP4Py runtime, returns a different iterator for each thread, which behaves according to the selected scheduling policy by returning the values assigned to that thread one by one. Once the iterator runs out of elements, it blocks until all other iterators have finished. This synchronization is achieved because all threads share a barrier object: once an iterator completes, it waits at the barrier until all threads reach it. As previously mentioned, this behavior can be avoided using the nowait clause.
Furthermore, the for directive can be extended to multiple nested loops using the collapse clause. Parallelizing multiple nested loops increases the number of iterations available for distribution among threads. This is especially useful when the outer loops have relatively few iterations but the total number of iterations across all loops is large. When the collapse clause is used, the OMP4Py runtime effectively flattens the nested loops into one loop. The number of nested loops to collapse is specified as an argument to the clause. The loops must be perfectly nested, and the number of iterations in the inner loops must not depend on the iteration variable of an outer loop. Figure 4 shows an example of using the collapse clause and its implementation process. The translation is similar to the previous case: the ranges of the two loops (lines 9-10, top code) are combined into a single loop where _omp_range (line 10, bottom code) receives the arguments of both ranges in tuple format and returns the values for both index variables. Finally, the example also illustrates the use of the lastprivate clause, which is similar to the private clause, but with the added feature that the variable is updated with the value assigned in the last iteration of the loop. A new variable (line 9, bottom code) is created to replace the original variable (line 6, top code) using the _omp_lastprivate function (line 14, bottom code), which employs an if statement to determine which thread performed the last iteration and updates the value of the original variable (line 15, bottom code).
1from omp4py import *
2
3@omp
4def f():
5 xs = [0] * 20
6 x = 0
7 with omp("parallel"):
8 with omp("for schedule(static, 2) collapse(2) lastprivate(x)"):
9 for i in range(len(xs)):
10 for i in range(4):
11 x = j
12 xs[i] += x
13 print(x, xs)
1from omp4py import *
2
3def f():
4 xs = [0] * 20
5 x = 0
6 def _omp_parallel_1():
7 nonlocal xs
8 nonlocal x
9 omp_x = None
10 for i, j in _omp_range((0, 0), (len(xs), 4), (1, 1), schedule=’static’, chunks=2):
11 omp_x_1 = j
12 xs[i] += omp_x_1
13
14 if _omp_lastprivate(i, j):
15 x = omp_x_1
16
17 _omp_parallel_run(_omp_parallel_1)
18
19 print(x, xs)
Figure 4: Example of the for directive with collapse and lastprivate clauses: user code (top) and its corresponding translation by OMP4Py (bottom).
3.2.2 Sections
The sections directive divides the work among threads by assigning each thread a structured block defined with the section directive. Therefore, only blocks with the section directive can exist within a sections directive. Each block will be executed once by a single thread, but the order of execution is not predetermined, and any thread may execute one or multiple blocks depending on availability. In sequential execution, the blocks are executed in the order of their definition. Similar to the for directive, there is a synchronization barrier that blocks all threads until all section blocks have been executed. This barrier can also be removed using the nowait clause.
Figure 5 shows an example of the sections directive with three section blocks (top) and its implementation by OMP4Py (bottom). Each block performs a print operation to display a number. In sequential execution, the output sequence is always ’1 2 3’, but with parallel execution, the numbers can be printed in any order. The implementation uses a with block along with the _omp_sections function (line 5, bottom code), which serves two purposes: first, it sets each section block as unexecuted, and second, it provides an exit barrier when all blocks have been executed. This is possible because the with block calls the __enter__ method on entry and the __exit__ method on exit. This mechanism ensures that if nowait is not used, a call to the barrier is made at the end of the block. Additionally, a unique identifier is assigned to each section block. The _omp_section function (lines 6, 8 and 10, bottom code) and the if block check whether the block has been executed or needs to be executed. If the block has not yet been executed, the function returns true and allows one thread to execute it. Afterward, any subsequent call to _omp_section for that block will return false, preventing additional threads from executing it.
Finally, although not shown in the example, the sections directive also supports the lastprivate clause, similar to the for directive detailed previously. Its internal implementation mirrors that of the for directive, using the identifier number of the last section block as an argument to _omp_lastprivate, which checks if the invoking thread executed that block.
1from omp4py import *
2
3@omp
4def f():
5 with omp("parallel"):
6 with omp("sections"):
7 with omp("section"):
8 print(1)
9 with omp("section"):
10 print(2)
11 with omp("section"):
12 print(3)
1from omp4py import *
2
3def f():
4 def _omp_parallel_1():
5 with _omp_sections():
6 if _omp_section(0):
7 print(1)
8 if _omp_section(1):
9 print(2)
10 if _omp_section(2):
11 print(3)
12
13 _omp_parallel_run(_omp_parallel_1)
Figure 5: Example of the sections directive: user code (top) and its corresponding translation by OMP4Py (bottom).
3.2.3 Single
The single directive specifies a block within a parallel region that will be executed only once by a single thread. The thread that executes the block is indeterminate; it will be the first thread to reach the block. In sequential execution, the directive has no effect since there are no additional threads to execute the block. As with other worksharing directives, the remaining threads must wait until the block’s execution is complete before proceeding. This barrier can also be removed using the nowait clause.
1from omp4py import *
2
3@omp
4def f():
5 x = 0
6 with omp("parallel firstprivate(x)"):
7 with omp("single copyprivate(x)"):
8 x += 1
9 print(x)
1from omp4py import *
2
3def f():
4 x = 0
5 def _omp_parallel_1():
6 _omp_x_1 = x
7 with _omp_single() as _omp_1:
8 if _omp_1:
9 _omp_x_1 += 1
10 _omp_copyprivate_set(_omp_x_1)
11 _omp_x_1 = _omp_copyprivate_get()
12
13 _omp_parallel_run(_omp_parallel_1)
Figure 6: Example of the single directive: user code (top) and its corresponding translation by OMP4Py (bottom).
Figure 6 shows an example of user code when using a single directive (top) and its internal implementation by OMP4Py (bottom). In this code, a variable declared with the firstprivate clause (line 6, top code) creates a private copy for each thread, initialized with the value from before the parallel region begins. Inside the single block, the variable is incremented by one (line 8, top code). After the single block finishes, the value of the variable from the thread that executed the block is broadcast to all other threads using the copyprivate clause. Note that variables listed in the copyprivate clause must be private to each thread. Once the broadcast is complete, all threads have the updated value and print it (line 9, top code).
The implementation of the single directive is similar to that of the sections directive, requiring a with block to control thread entry and exit. The _omp_single function (line 7, bottom code) returns true only for the first thread to invoke it, returning false for all other threads. This return value is stored in a temporary variable, and the if block ensures that only the thread that received true executes the directive’s code (line 8, bottom code). The copyprivate clause adds a call to the _omp_copyprivate_set function at the end of the block (line 10, bottom code), which takes all variables listed in the clause as arguments. These values are stored internally and then updated in other threads using _omp_copyprivate_get (line 11, bottom code) through multiple assignments. The synchronization barrier in the with block ensures that no thread calls _omp_copyprivate_get before _omp_copyprivate_set has been invoked by the thread that executed the directive. Removing this barrier with nowait would require a more complex, inefficient, and hard-to-debug mechanism. Consequently, the OpenMP standard restricts the use of copyprivate and nowait together in the same single directive.
3.3 Tasking Directives
OpenMP tasking provides a flexible way to parallelize heterogeneous or dynamic workloads. The task directive allows for the creation of tasks, which are units of work that can be executed by any thread in the team. This is particularly useful for applications with recursive algorithms, irregular loops, or other patterns where the workload cannot be distributed uniformly across threads from the beginning. When a thread reaches a task directive, it wraps the associated block of code and its data environment into a task, which can then be executed by any thread in the current team. A task does not need to be executed immediately, it can be stored in a queue for later execution. Threads in the team can dynamically take tasks from the queue and execute them. Tasks can be executed explicitly with the taskwait directive or implicitly when a thread becomes available. In any case, all tasks will be completed before the thread team terminates.
The implementation of the task directive is similar to parallel in the sense that parallel can be considered as a task executed by all threads in the team at its start. The definition of the variable scope for creating the environment follows the same principles, utilizing clauses such as default, private, firstprivate, and shared. There is an additional if clause that allows for conditional task creation: if the condition evaluates to false, the task is executed immediately instead of being enqueued.
1from omp4py import *
2
3@omp
4def fib(n):
5 i = 0
6 j = 0
7 if n < 2:
8 return n
9 with omp("task"):
10 i = fib(n - 1)
11 with omp("task"):
12 j = fib(n - 2)
13 omp("taskwait")
14 return i + j
15
16@omp
17def f(n):
18 x = 0
19 with omp("parallel"):
20 with omp("single"):
21 x = fib(n)
22 print(x)
1from omp4py import *
2
3def fib(n):
4 i = 0
5 j = 0
6 if n < 2:
7 return n
8 def _omp_task_1():
9 nonlocal i
10 nonlocal j
11 i = fib(n - 1)
12 _omp_task_submit(_omp_task_1)
13 def _omp_task_2():
14 nonlocal i
15 nonlocal j
16 j = fib(n - 2)
17 _omp_task_submit(_omp_task_2)
18 _omp_taskwait()
19 return i + j
20
21def f():
22 x = 0
23 def _omp_parallel_1():
24 nonlocal x
25 with _omp_single() as _omp_1:
26 if _omp_1:
27 x = fib(n)
28
29 _omp_parallel_run(_omp_parallel_1)
Figure 7: Fibonacci calculation example using the tasking directives: user code (top) and its corresponding translation by OMP4Py (bottom).
Figure 7 shows the recursive calculation of Fibonacci numbers using tasks (top) and its internal implementation generated by OMP4Py (bottom). The algorithm defines the Fibonacci function (line 4, top code), which calculates the value for a given $n$ by summing the recursive calls to the same function for the values $n-1$ and $n-2$ (lines 10 and 12, top code). The initial call to the Fibonacci function is placed in a single block to ensure it is executed by only one thread (line 21, top code), while the remainder threads are blocked. The function then creates a task for the recursive calculation (lines 9 and 11, top code) and blocks the current task until they have been executed (line 13, top code). The threads in the team will consume the created tasks and generate new tasks until they reach the base cases of $n=0$ and $n=1$. The code generated by OMP4Py shows the similarities with the parallel directive. In this way, the code within the task directive (lines 9 and 11, top code) and the parallel directive (line 19, top code) both result in the creation of functions: _omp_task_1, _omp_task_2 (lines 8 and 13 in the bottom code), and _omp_parallel_1 (line 23, bottom code), respectively. The only notable difference is that parallel uses the function _omp_parallel_run (line 29, bottom code), while task employs _omp_task_submit (lines 12 and 17, bottom code) which places the task function into a shared queue. The queue is accessible by all threads that will be able to pick up the task and execute it. Finally, the taskwait clause is implemented as a simple function (line 18, bottom code), which forces the caller thread to consume all tasks in the queue until it is empty, at which point the function returns.
3.4 Interaction with Python
In this section, we will explain the implementation of the _omp internal OMP4Py API functions that were introduced in the previous sections and how they interact with Python. The generated code uses the internal API to manage operations such as thread management and work distribution, all implemented with Python code and standard libraries, making OMP4Py a pure native Python library.
When a Python program using OMP4Py is executed, a single thread begins by running the main program, as in a typical Python script. However, it is important to note that before the main program starts, the Python interpreter processes all decorators, including omp decorators. So, by the time the main program starts, all OpenMP directives within the source code will be replaced with a new pure Python version with parallel behavior. According to the OpenMP standard, this initial thread is called the master thread. OMP4Py initializes the master thread context with the first call to any function in its API. The remaining threads created in parallel constructs will be initialized with a context derived from the master thread before they begin execution. For this reason, any thread created outside the OMP4Py clauses, for instance, using asyncio, concurrent.futures, threading, or the multiprocessing module, will not have an associated context and will be treated by OMP4Py as an uninitialized master thread. If there is a subsequent call to any function in the OMP4Py API by one of these threads, a new OMP4Py context will be created for it. In this way, the thread will act as a new master thread and can independently use all the features of OMP4Py, leaving the user responsible for managing potential concurrency issues between different instances.
The context in OMP4Py is implemented by a stack object where OpenMP construct tasks are stored. In the OpenMP standard, the master thread exists within a single-threaded parallel region. Therefore, the initialization of the master thread’s context involves adding a parallel construct task to the stack, allowing the OpenMP API (like omp_get_thread_num) to be called from any point in the source code, ensuring consistent results. The context is stored locally to the thread using threading.local.
In a typical execution flow of an OpenMP application, the next task added to the stack is the parallel task, which creates a parallel region with multiple threads. The _omp_parallel_run function is responsible for this process. It works as follows: the master thread creates a context for each new thread, with a stack that is a shadow copy of its own. The master thread then adds the parallel task to all stacks, including its own, and finally, using the threading library, the threads are created with their respective contexts as arguments. The parallel task contains the following information: the thread ID, the number of threads, a mutex, a barrier, a shared task list, and a shared dictionary. The thread ID and number of threads are critical for work distribution and API functions. The mutex and barrier are used for constructs like critical and barrier. The shared task list holds tasks (created by the task construct) that must be completed before the parallel region ends, while the shared dictionary stores common data, such as iteration counters for dynamic workloads in a for construct. Exceptions thrown within a parallel region are not automatically propagated to the master thread. Instead, the behavior depends on how the exception is handled within each thread: if an exception occurs, it must be caught and handled within that thread. OMP4Py catches unhandled exceptions in parallel regions to prevent the program from terminating unexpectedly, but relying on this behavior is considered bad practice.
As explained in Section 3.2.1, the for construct is executed by each thread in the parallel region using the _omp_range function, which returns an iterator for the thread’s assigned values. For static scheduling, each thread divides the work based on its thread ID and returns an independent iterator. For dynamic or guided scheduling, the thread acquires the mutex, checks the shared dictionary for an iteration counter (creating it if needed), and uses it to fetch new work chunks. This is done using Python iterators and the yield function. The sections and section constructs (Section 3.2.2) work similarly, where _omp_sections creates a sections task with an empty set, and each _omp_section function checks and updates the set to track executed sections, all while using the mutex to avoid race conditions.
The _omp_single function (see Section 3.2.3) ensures that only one thread within a parallel region executes a section of code. It does so by returning the mutex of the region, which locks the code block for exclusive access by one thread using the with statement. This causes the mutex to be locked at the start of the block and released automatically when the block finishes.
The functions _omp_copyprivate_get and _omp_copyprivate_set facilitate data sharing between threads. _omp_copyprivate_get retrieves variables from the shared dictionary and makes threads wait for data using threading.Condition, while _omp_copyprivate_set stores data and notifies other threads when it becomes available.
Finally, the _omp_task_submit function (Section 3.3) adds a task to the shared task list of the current parallel region, enabling it to be queued for execution. Tasks are then assigned to available threads for execution, either when the parallel block finishes or through the taskwait construct. The _omp_taskwait function ensures that a thread consumes tasks from the shared list, only returning when the list is empty. To prevent race conditions, a mutex is used each time a thread removes a task for execution.
4 Experimental Results
Next, we will evaluate the performance and scalability of different Python applications parallelized using OMP4Py. Additionally, we will highlight the differences and advantages of having a native OpenMP implementation in Python, as provided by OMP4Py, compared to PyOMP, which is a Numba-based prototype with OpenMP support.
Experiments were conducted using one server with two 24-core Intel Xeon Gold 5220R @2.2GHz processors and 192 GB of RAM. The software used was Python v3.13, NumPy v2.1.1, mpi4py v4.0.0, NetworkX v3.3, and PyOMP v0.1 (September 2024). Execution times were averaged over 10 measurements for each test.
Figure 8: Scalability of the different parallel numerical applications. Axis in log scale.
4.1 Numerical algorithms
We have selected six algorithms that represent different types of numerical application patterns to evaluate the performance and scalability of OMP4Py. In particular:
  • 1.
    Fast Fourier Transform (fft). It is an efficient algorithm used to compute the Discrete Fourier Transform (DFT) of a sequence, allowing for the conversion of a signal from its time domain to its frequency domain. Performance tests were run using a complex data vector of 4 million numbers.
  • 2.
    Jacobi method (jacobi). It is an iterative algorithm for solving systems of linear equations of the form $A\cdot x=b$, where $A$ is a matrix, and $x$ and $b$ are vectors. At each iteration, the solution vector is updated based only on values from the previous iteration. We used a square matrix $A$ of size $1k\times 1k$, performing up to 1,000 iterations, with a stopping criterion of an error tolerance of $1\times 10^{-6}$.
  • 3.
    LU decomposition (lu). It is a method used to factor a matrix $A$ into the product of two matrices: a lower triangular matrix $L$ and an upper triangular matrix $U$, such that $A=L\cdot U$. This factorization simplifies solving systems of linear equations, matrix inversion, and determinant computation. We applied LU decomposition to a square matrix of size $1k\times 1k$.
  • 4.
    Molecular dynamics simulation (md). The simulation was conducted to study the motion of particles over time. The velocity Verlet integration scheme was employed to update positions, velocities, and accelerations. We simulated a system of 2,000 particles interacting with a central pair potential.
  • 5.
    Computing $\pi$ (pi). The area under the curve $y=\frac{4}{1+x^{2}}$ between 0 and 1 provides an approximation for $\pi$. This integral can be estimated using numerical summation, where we used 2 billions of intervals to compute the approximation.
  • 6.
    QUAD. It is a numerical integration technique that estimates the value of an integral using an averaging method. It approximates the integral of the function $f(x)=\frac{50}{\pi\cdot(2500\cdot x^{2}+1)}$ over the interval from $A=0$ to $B=10$. This method involves sampling the function at numerous points within the interval to compute an average value, which is then used to estimate the integral. For our tests, we employed 1 billion iterations.
All the source codes can be found in the OMP4Py repository. Figure 8 shows the execution times of the different parallel benchmarks using from 1 to 24 threads. It can be observed that the scalability of the OMP4Py applications is not good. For instance, only 4 out of 6 benchmarks benefit from using 4 threads, while using more threads consistently proves detrimental to performance. The best speedup overall is 3.18$\times$, obtained for the molecular dynamics application and 4 threads. It is important to highlight that these scalability problems are not related to the Python code generated by OMP4Py. As we will show in Section 4.2, non-numerical OMP4Py applications exhibit strong scalability with codes containing the same structure and OpenMP directives as those discussed in this section. Therefore, these results demonstrate that the current version of the Python interpreter (v3.13) still lacks mature support for multithreading, with several open implementation issues that need to be addressed to exploit it efficiently. While this new version is an important step forward in removing the GIL, it still limits the use of OMP4Py to a small number of threads when running numerical applications.
Finally, for illustrative purposes, Figure 8 also shows the performance results obtained by the applications when compiled using PyOMP (Numba). As expected, they are significantly faster than pure Python, as Numba compiles Python code into optimized machine code, bypassing the slower interpreted execution of standard Python. This leads to substantial reductions in the execution times, particularly for compute-intensive tasks such as numerical computations. In any case, we have observed problems in the scalability of some benchmarks (see fft and md results). Additionally, PyOMP was unable to execute the jacobi method using more than one thread.
4.2 Applications using non-numerical libraries and Python objects
As we pointed out previously, PyOMP is a fork of the Numba project. The Numba library’s njit decorator accelerates Python functions by compiling them to machine code, but it imposes limitations on using features such as functions from non-Numba-optimized libraries and certain Python objects and data structures. Numba provides special support for some commonly used libraries like math and numpy, allowing their functions to be called within @njit-compiled functions, as these libraries are optimized for machine code compilation. As a result, Numba restricts function usage to those either decorated with @njit or from libraries that Numba specifically optimizes. In contrast, OMP4Py, being a native OpenMP implementation for Python, is designed to work with a broader range of code, including those that may not be compatible with Numba’s restrictions.
To illustrate the benefits of having this native OpenMP implementation for Python, we have implemented the following applications using OMP4Py:
  • 1.
    Graph Clustering. The clustering coefficient of a node is the fraction of possible triangles that pass through that node in an unweighted graph. We used a graph with 300k vertices, each connected by 100 edges, as input. The graph generation, storage, and clustering algorithm were implemented using the NetworkX [15] library. Note that PyOMP cannot run this benchmark because Numba is unable to compile the object Graph and the clustering algorithm function calls, as they are part of an external library that is not optimized for Numba.
  • 2.
    Wordcount. It is a simple algorithm that counts the number of occurrences of each word in an input text. We generated a text consisting of 1 million characters, featuring words with lengths between 3 and 10 letters, with a 10% probability that a new line will be added after each word. Note that although more recent versions of Numba have increasingly added experimental support for Python dictionaries, PyOMP is a fork of an earlier version that lacks the necessary support to compile Wordcount dictionaries.
Source codes can be found in the OMP4Py repository.
Figure 9: Scalability of the Graph Clustering (left) and Wordcount (right) applications. Axis in log scale.
Figure 9 shows the scalability of the Graph Clustering and Wordcount applications using up to 48 threads. The scalability of OMP4Py is good, especially for Wordcount, which achieves a 25.5$\times$ speedup compared to the sequential execution when using 48 threads. This behavior is very different from the performance results obtained for the numerical applications (see Figure 8). However, the structure and OpenMP directives used in both types of applications are very similar, with the only differences being related to the operations performed inside the loops. For example, when closely examining the OMP4Py codes in Figure 10 for the $\pi$ calculation (left) and the Wordcount application (right), it can be observed that the OpenMP clauses are identical, with only the operations in the $i$ loop and the critical block differing. As a result, the scalability issues detected for the $\pi$ calculation cannot be attributed to OMP4Py, since the code generated for both applications is essentially the same. Therefore, these results can only be explained by the fact that free-threaded support for Python 3.13 is still in an experimental stage, leaving plenty of room for improving performance and efficiency.
1from omp4py import *
2
3@omp
4def pi(n):
5 w = 1.0 / n
6 PI = 0.0
7 with omp("parallel"):
8 pi_local = 0.0
9 with omp("for"):
10 for i in range(n):
11 local = (i + 0.5) * w
12 pi_local += 4.0 / (1.0 + local * local)
13 with omp("critical"):
14 PI += pi_local
15
16 return PI * w
1from omp4py import *
2
3@omp
4def wordcount(lines):
5 count = {}
6 with omp("parallel"):
7 local_count = {}
8 with omp("for"):
9 for i in range(len(lines)):
10 for word in lines.split():
11 if word in local_count:
12 local_count[word] += 1
13 else:
14 local_count[word] = 1
15 with omp("critical"):
16 count.update(local_count)
17
18 return count
Figure 10: OMP4Py user codes for the Pi (left) and Wordcount (right) applications.
4.3 Hybrid applications combining OMP4Py with mpi4py
One of the main features of OMP4Py is that it can be combined with mpi4py [4] to implement hybrid parallel applications that can exploit intra- and inter-node parallelism. The mpi4py package provides Python bindings for the Message Passing Interface (MPI) standard [16], which is the most widely used and dominant programming model in HPC.
It is important to highlight that although mpi4py is a Python wrapper for the MPI C library, Numba cannot use MPI code within its functions because it treats mpi4py as an external library. Numba’s compilation process focuses on translating Python code to machine code but does not include functionality for integrating or compiling external libraries like MPI automatically. Even though mpi4py interfaces with C, Numba is unaware of how to compile MPI operations into their C equivalents. For this reason PyOMP cannot be combined with mpi4py.
As a case study, we implemented a hybrid parallel version of the Jacobi application described in Section 4.1. To implement the Jacobi method to solve $Ax=b$ using MPI, the matrix $A$ and vector $b$ are distributed across multiple processors, with each processor managing a portion of the matrix rows and corresponding elements of the vector. During each iteration, processors compute updated values of $x$ using OpenMP based on their assigned subset of $A$ and $b$. MPI function MPI_Allgather is employed to exchange the updated vector $x$ among all processors to ensure consistency. Convergence is assessed by computing the global error, with MPI_Allreduce used to aggregate and verify whether the stopping criterion is satisfied. This code can also be found in the OMP4Py repository.
Figure 11: Scalability of a hybrid implementation (OMP4Py + mpi4py) of the Jacobi application using different number of computing nodes. 16 threads per computing node were used. Axis in log scale.
Experiments in this section were performed on a cluster using up to 8 computing nodes, each node containing two 24-core Intel Xeon Gold 5220R @2.2GHz processors and 192 GB of RAM. The performance results, illustrated in Figure 11, show speedups of 1.84$\times$, 3.13$\times$ and 4.85$\times$ when using 2, 4, and 8 nodes respectively, compared to using a single node. These strong scalability results demonstrate the potential of OMP4Py to enhance the performance of applications currently parallelized using only mpi4py. Tests were conducted using 16 threads per node.
5 Conclusions
In this paper, we introduce OMP4Py [^4] It is publicly available at https://github.com/citiususc/omp4py, a novel implementation of the OpenMP standard designed specifically for Python. It fully supports the API specifications of version 3.0. OMP4Py integrates OpenMP into Python by adapting its directive-based approach, allowing users to embed parallel constructs directly into their code. These transformer directives enable parallel execution by instructing OMP4Py to modify the Python code as needed.
The performance evaluation of OMP4Py reveals limited scalability when using numerical algorithms. Specifically, only 4 out of 6 benchmarks showed improvement with 4 threads, with performance typically degrading when more threads were used. However, we demonstrated that these scalability issues are due to Python’s multithreading limitations in version 3.13, rather than OMP4Py’s implementation.
Non-numerical applications using Python objects and external libraries show much better scalability, achieving up to 25.5$\times$ speedup with 48 threads. This confirms that OMP4Py works well for a broader range of applications compared to PyOMP (Numba), which is restricted to numerical algorithms by its inability to handle these types of data structures and libraries.
Finally, OMP4Py can be effectively combined with mpi4py to create hybrid parallel applications that leverage both intra- and inter-node parallelism, demonstrating strong scalability.
In the future, OMP4Py will be extended to support newer versions of the OpenMP standard, including versions 4.0 through 6.0. This will involve incorporating advanced features such as task dependencies and thread teams. Additionally, the current implementation will be optimized by reducing mutex locks and exploring atomic operations, which are commonly used in C-based OpenMP implementations.